### v1

# AgentExpt: Automating AI Experiment Design with LLM-based Resource Retrieval Agent

Abstract.
Large language model agents are becoming increasingly capable at web-centric tasks such as information retrieval, complex reasoning and online navigation. These emerging capabilities have given rise to surge research interests in developing LLM agents for facilitating scientific quest, which is often known as AI Scientist systems. One key application in AI research is to automate AI experiment design through agentic dataset and baseline retrieval, which have the potential to largely accelerate AI experiments and assist peer review. However, prior efforts suffer from limited data coverage, as recommendation datasets primarily harvest candidates from public portals and omit many datasets actually used in published papers, and from an overreliance on content similarity that biases models toward superficial similarity and overlooks experimental suitability. Harnessing collective perception embedded in the baseline and dataset citation network, we present a comprehensive framework for baseline and dataset recommendation. First, we design an automated data-collection pipeline that links roughly one hundred thousand accepted papers to the baselines and datasets they actually used. Second, we propose a collective perception enhanced retriever. To represent the position of each dataset or baseline within the scholarly network, it concatenates self-descriptions with aggregated citation contexts. To achieve efficient candidate recall, we finetune an embedding model on these representations. Finally, we develop a reasoning-augmented reranker that exacts interaction chains to construct explicit reasoning chains and finetunes a large language model to produce interpretable justifications and refined rankings. The dataset we curated covers 85% of the datasets and baselines used at top AI conferences over the past five years. On our dataset, the proposed method outperforms the strongest prior baseline with average gains of +5.85% in Recall@20, +8.30% in HitRate@5, and +7.90% in HitRate@10. Taken together, our results advance reliable, interpretable automation of experimental design.
baseline recommendation, dataset recommendation, citation contexts, reranking, LLM agents
Figure 1. Overview of the Research Problem:
1. Introduction
Recent advances in large language models (LLMs) have rapidly expanded their capabilities(Brown et al., 2020; Radford et al., 2019, 2018; Shang et al., 2024b, a; Li et al., 2025). In particular, LLM-powered agents have been widely applied to web-centric tasks, including web navigation and browsing (Zhou et al., 2023; Deng et al., 2023; Chezelles et al., 2024; Lù et al., 2024), personalized recommendation and user modeling (Zhang et al., 2024; Wang et al., 2023; Hao et al., 2025), and goal-oriented online shopping (Yao et al., 2022). At the same time, there is growing interest in leveraging LLM agents to gather information from the web for scientific tasks such as literature review and idea generation (Huang et al., 2025; Team et al., 2025). Building on this trend, end-to-end research frameworks such as Carl (Institute, 2025), Zochi (Intology, 2025), and the AI Scientist (Lu et al., 2024; Yamada et al., 2025) have been proposed to automate the research pipeline. However, these systems typically operate within narrow, pre-defined problem settings and do not require careful selection of baselines or datasets, which limits their generality and rigor. Meanwhile, the number of available models and datasets in machine learning has grown dramatically, making it increasingly important and challenging to recommend appropriate baselines and datasets for a new research idea, see fig 1.
Figure 2. Constructing chain-derived candidates and analysis. Left: From the interaction graph of papers, baselines and datasets, we extract interaction chains and aggregate the terminal items to form a chain-derived dataset/baseline pool. Right: (i) Recall (%) between a target paper’s actual baselines/datasets and candidates from each setting, and (ii) Precision (%) of overlapped items within the corresponding candidate pool. We evaluate three settings on both the baseline side and the dataset side: chain-derived top–100, same conference-derived top–100, embedding top- 100. Chain-derived candidates recover on average 60.14% of baselines and 78.61% of datasets while occupying 2.52% and 5.63% of the respective chain-derived pools, indicating that interaction chains provide a compact yet highly informative prior for selecting baselines and datasets.:
Research on dataset and baseline recommendation spans two components: datasets and methods. On the dataset side, RecBaselines2023 targets baseline recommendation within recommender systems (Ivanova et al., 2023), while DataFinder and DataHunter address dataset recommendation from problem statements or open retrieval (Viswanathan et al., 2023a; Färber and Leisinger, 2021b, c). On the method side, representative approaches include supervised text classification (Färber and Leisinger, 2021c, b), dense bi encoder retrieval (Viswanathan et al., 2023a), graph based modeling over papers and datasets (Altaf et al., 2019; Qayyum et al., 2025), and collaborative filtering over baseline co occurrences (Ivanova et al., 2023). Collectively, these strands establish initial tasks and evaluation setups. Despite progress, important gaps remain. On the data side, existing baseline recommendation dataset is limited to specific domains (Ivanova et al., 2023) and scales, while dataset recommendation benchmarks primarily harvest candidates from portals such as Papers with Code (Viswanathan et al., 2023a) or Wikidata (Färber and Leisinger, 2021c, b),omitting many datasets that are actually used in published papers. This limits coverage and reduces practical utility for real research workflows. On the method side, most systems (Färber and Leisinger, 2021b; Beltagy et al., 2019; Qayyum et al., 2025; Goyal et al., 2024; Gu et al., 2022) rely solely on first-person self-descriptions which biases models toward surface-level semantic similarity and yields items that appear topically related yet are ill-suited to the intended experimental setting. Moreover, prior work typically treats baseline and dataset recommendation independently, overlooking their synergy.
Motivated by our day-to-day research practice, we typically begin from a set of closely related papers and follow their trails to the baselines and datasets they actually used, downloading artifacts directly rather than browsing portal listings. In choosing these components, we consider baselines and datasets jointly, since pairings that have been used together—or are methodologically compatible—are more likely to fit our intended evaluation setting. Moreover, when writing a paper we must justify these choices, situating a baseline or dataset relative to prior work. So the citation contexts constitute third-person signals that complement an asset’s first-person self-description. These observations motivate a unified formulation that links papers, baselines, and datasets and explicitly leverages community-level perception.
Building on this motivation, to build a massive baseline and dataset recommendation dataset that explicitly links research papers to the baselines and datasets they actually use, we design an automated data-collection pipeline that gathers roughly one hundred thousand accepted papers from the past decade across ten flagship AI venues. We convert PDFs to structured text, identify and parse experiment-centric sections, and apply LLM-assisted extraction to recognize true mentions of baseline methods and evaluation datasets. We then normalize naming variants, merge duplicates, and associate each entity with its introducing paper, yielding a high-precision mapping from papers to the experimental components they employ. On this dataset, to capture how the community actually positions and uses a baseline or dataset, we extract citation contexts as third-person perception signals that complement first-person self-descriptions. For each target (baseline or dataset), we summarize its citation contexts and represent it by concatenating its self-description with the aggregate community perception. To efficiently recall relevant candidates, we finetune an embedding model on these representations. To exploit baseline–dataset synergy(see fig 2) and refine the shortlist, we develop a reasoning-augmented reranking framework that extracts paper–baseline–paper-dataset and paper–dataset–paper-baseline interaction chains and uses them to construct reasoning chains. To stimulate reasoning ability, We finetune a language model to generate these chains for reranking, producing interpretable judgments that leverage the interaction-chain prior and go beyond simple semantic matching. This two-stage pipeline couples collective perception–enhanced retrieval with interaction chain–driven reasoning, yielding recommendations that are both semantically grounded and aligned with evaluation intent.
We validate the proposed framework on AgentExpt, our new corpus linking papers (2015–2025, ten flagship AI venues) to baselines and datasets. We evaluate two tasks—baseline and dataset recommendation—under Recall@$k$ and HitRate@$k$. Compared with the strongest prior baseline , our method achieves averaged improvements across the two tasks of +5.85% in Recall@20, +8.30% in HitRate@5, and +7.90% in HitRate@10. Ablations show that collective perception contributes the largest share of retrieval gains, while interaction-chain reasoning drive consistent improvements in hitrate.
The key contributions of this work are as follows:
  • We construct, to our knowledge, the first large-scale, high-quality dataset for baseline and dataset recommendation. Our dataset covers roughly 85% of the baselines and datasets used at top AI conferences over the past five years.
  • We propose a collective perception–enhanced retriever that represents each candidate by jointly encoding its self-description and the community’s perception of how it is positioned and used, yielding retrieval signals that go beyond surface-level textual similarity and better reflect the evaluation intent.
  • We develop a reasoning-augmented reranking framework that leverages interaction chains to construct explicit reasoning chains and finetunes a large language model to produce interpretable justifications and refined rankings.
  • We empirically demonstrate state-of-the-art performance on both baseline and dataset recommendation: averaged over the two tasks, HitRate@5 improves by 8.30%, HitRate@10 by 7.90%, and Recall@20 by 5.85% over the strongest prior baseline.
2. Related Work
2.1. LLM-Based Agents for Research Assistance
Recent efforts have explored the use of LLM-based agents to automate scientific research, ranging from idea generation to full paper drafting. Systems such as ResearchAgent (Baek et al., 2024), IdeaSynth (Pu et al., 2025), and NovelSeek (Team et al., 2025) focus on generating and refining research ideas with literature grounding and structured feedback. In parallel, a line of research targets algorithm optimization itself, where agents search for improved procedures or code, including AlphaEvolve (Novikov et al., 2025), FunSearch (Fawzi and Paredes, 2023). Building on this, end-to-end frameworks including The AI Scientist (Lu et al., 2024; Yamada et al., 2025), Zochi (Intology, 2025), Carl (Institute, 2025), AI-Researcher (Tang et al., 2025), Agent Laboratory (Schmidgall et al., 2025), and domain-specific agents such as Biomni (Huang et al., 2025) attempt to close the entire loop by combining literature review, idea generation, and paper writing. While these systems represent significant progress toward autonomous scientific discovery, they overlook a crucial step in practical research, namely the selection of appropriate baselines and datasets.
2.2. Benchmarks for Baseline and Dataset Recommendation
Early work relied on general dataset search engines that index web metadata and support keyword queries, but they lack task-specific recommendation benchmarks tied to research problem descriptions or scientific contexts (Brickley et al., 2019; Chapman et al., 2020). For baseline recommendation, RecBaselines2023 is, to our knowledge, the only dataset that recommends baselines conditioned on a paper, but its scope is confined to the recommender systems domain and its scale is limited to on the order of several hundred papers and baselines (Ivanova et al., 2023). For dataset recommendation, Färber and Leisinger (2021c) cast the task as multi label classification from natural language problem statements to a fixed inventory of datasets, which simplifies evaluation but cannot naturally handle unseen datasets at test time; Viswanathan et al. (2023a) instead formulate an open domain retrieval benchmark with a large automatically constructed training set and a smaller expert annotated test set, showing that bi encoder retrieval substantially outperforms keyword search and third party engines. However, both lines of work primarily source candidate datasets from portals such as Papers with Code, OpenAIRE, and Wikidata, which means many datasets actually used in published top venue papers but not indexed by these portals are omitted, reducing coverage and limiting practical utility for real research workflows.
2.3. Methods for Baseline and Dataset Recommendation
Prior work spans three main paradigms. A first line treats dataset recommendation as supervised text classification from problem descriptions to a fixed label set of datasets, which is simple and effective on closed inventories but struggles with unseen labels and label drift at test time (Färber and Leisinger, 2021c, b). A second line builds heterogeneous graphs over papers, methods, and datasets and learns links for recommendation, for example via variational graph autoencoders; these approaches leverage co usage and citation topology but face cold start issues and are less natural for free form queries (Qayyum et al., 2025; Altaf et al., 2019). A third line adopts neural retrieval with dense encoders, where a bi encoder retrieves datasets for natural language ideas at open world scale and typically outperforms keyword search, though most systems optimize semantic similarity rather than experimental suitability and offer limited explainability (Viswanathan et al., 2023a). For baseline recommendation, collaborative filtering over paper baseline co occurrences has been shown effective but is domain bound and provides little semantic justification for new ideas (Ivanova et al., 2023).
Figure 3. Pipeline for constructing the AgentExpt knowledge base. We (1) download and parse papers (from flagship AI conferences), (2) identify baselines and datasets by locating experiment sections, (3) apply rule-based and LLM-based filtering using citation frequency, naming consistency, and contextual positioning to prune false positives. The final dataset contains 108,825 papers, 116,970 baseline entities, 68,316 dataset entities, and their respective cross‑entity connections.:
3. Preliminary
We formalize baseline and dataset recommendation as a retrieval task in the context of scientific research. Given a research idea expressed as a natural language query $q$, the objective is to retrieve from a large candidate pool the most suitable subset of experimental components, including baselines and datasets, that are appropriate for evaluating the proposed idea.
Table 1. Comparison of datasets for recommending baselines or datasets.
Dataset # Papers # Baselines # Datasets Top conf Repo links
AgentExpt 108,825 116,970 68,316
DataFinder 17,495 N/A 7,000
RD4SPD 88,000 N/A 1,413
RecBaselines 903 363 N/A
3.1. Problem Formulation
Let $\mathcal{Q}$ denote the space of research ideas, $\mathcal{B}$ the inventory of available baselines, and $\mathcal{D}$ the inventory of datasets. For a given query $q\in\mathcal{Q}$, the system should return two subsets:
$R_{B}(q)\subseteq\mathcal{B},\quad R_{D}(q)\subseteq\mathcal{D},$
where $R_{B}(q)$ and $R_{D}(q)$ represent the recommended baselines and datasets respectively. These outputs are intended to approximate the ground-truth relevance sets
$G_{B}(q),\quad G_{D}(q),$
Unlike generic document retrieval, this task requires reasoning not only about topical similarity but also about experimental suitability.
3.2. Input Representation
We instantiate the query $q$ using the abstract summary of a paper that expresses the research idea. Formally, for a paper $p$ with abstract $a(p)$, we set
$q\;=\;\mathrm{Summ}\!\big(a(p)\big),$
where $\mathrm{Summ}(\cdot)$ is implemented with GPT4o (OpenAI, 2024) to produce a concise synopsis that preserves the problem setting, assumptions, and evaluation intent. This abstract based representation captures the key scientific context.
3.3. Evaluation Metrics
Following conventions in information retrieval, we evaluate recommendation performance by comparing the retrieved subsets $(R_{B}(q),R_{D}(q))$ against the gold sets $(G_{B}(q),G_{D}(q))$. The primary metrics are:
  • Recall@$k$: The fraction of gold items (baselines or datasets) that appear in the top-$k$ retrieved results. High recall ensures that critical experimental components are not missed.
  • HitRate@$k$: A binary measure that equals 1 if at least one relevant item is retrieved within the top-$k$, and 0 otherwise. This reflects a user-centric view of success: whether the system surfaces some useful component early.
4. Data Collection
4.1. Overview
To effectively support the retrieval and recommendation of state-of-the-art baselines and suitable evaluation datasets corresponding to specific research topics, it is imperative to precisely extract baselines and datasets utilized in empirical evaluations reported in scholarly literature. Accomplishing this task requires extensive extraction of structured entities accompanied by comprehensive metadata associated with respective papers, baselines, and datasets. However, current academic resources (e.g., Papers with Code(Pap, [n. d.]), Hugging Face(Hug, [n. d.])) and automated information extraction methodologies (e.g., DataFinder(Viswanathan et al., 2023b), DataHunter(Färber and Leisinger, 2021a)) exhibit substantial limitations concerning dataset coverage comprehensiveness, entity metadata completeness, and the accuracy of identifying PaperBaseline, PaperDataset associations.
Specifically, establishing a robust and high-quality Paper-Baseline-Dataset data involves addressing two primary challenges. First, significant heterogeneity in data sources and formats complicates information extraction efforts. Baselines often appear in abbreviated or inconsistent terminologies across different publications, while datasets frequently suffer from ambiguous naming conventions and inconsistent references. Moreover, the absence of explicit links within papers to original baseline implementations and dataset repositories further complicates automated text parsing, entity recognition, and information alignment processes. Second, the exponential growth of artificial intelligence research outputs exacerbates inefficiencies and scalability constraints inherent to manual or semi-automated annotation methods. Leading AI conferences receive tens of thousands of submissions annually, making traditional annotation approaches impractical and highlighting the need for scalable automation.
Motivated by recent advancements in automated scientific information extraction methodologies such as DataFinder (Viswanathan et al., 2023b), we have developed and implemented a comprehensive, end-to-end automated pipeline for scholarly data extraction and validation. Our pipeline specifically targets research papers published between 2015 and 2025 across ten prominent artificial intelligence conferences: NeurIPS (Neu, [n. d.]), ICML (ICM, [n. d.]), ICLR (ICL, [n. d.]), ACL (ACL, [n. d.]), NAACL (NAA, [n. d.]), EMNLP (EMN, [n. d.]), CVPR (CVP, [n. d.]), ICCV (ICC, [n. d.]), AAAI (AAA, [n. d.]), and IJCAI (IJC, [n. d.]). As a result, we have constructed a scholarly knowledge repository distinguished by extensive data coverage and remarkable accuracy.
4.2. Automated Data Collection Pipeline
The proposed automated pipeline comprises critical stages including paper acquisition and metadata extraction, structured content parsing, candidate entity recognition for baseline methods and datasets, and a rigorous validation process employing both rule-based filtering and large language model (LLM)-based verification. This pipeline significantly enhances entity extraction precision and recall. Furthermore, it systematically integrates essential metadata attributes for each entity, such as repository links, publication year, original data sources, evidence supporting entity recognition, and standardized canonical naming conventions. The end-to-end automation effectively mitigates challenges associated with data heterogeneity and substantially alleviates scalability issues encountered by manual annotation methodologies, thereby providing robust, accurate, and valuable data for academic recommendation and literature retrieval tasks.
Figure 4. The coverage of experiment baselines and datasets. The vertical axis indicates the coverage, calculated as the fraction of resources employed in year N that were introduced in preceding years (1 to N-1), reflecting the dependency on established experimental components over time.:
Acquisition and Parsing.
We systematically acquired research papers and corresponding metadata published between 2015 and 2025 from prominent AI academic repositories, including OpenReview (Ope, [n. d.]), arXiv (arX, [n. d.]), and official conference websites of ten leading AI venues. The collected metadata includes comprehensive information such as paper titles, authors, abstracts, publication years, conference affiliations, and direct links to publications. After rigorous deduplication and cleaning, the corpus exceeding one hundred thousand papers was refined into a high-quality dataset suitable for subsequent analyses. Utilizing the MinerU tool (Wang et al., 2024), we performed structured parsing on each paper, accurately extracting chapter structures, paragraph boundaries, table arrangements, and citation anchors, thereby ensuring reliable downstream entity extraction.
Baseline and Dataset Identification.
To identify baseline methods and datasets used in experiments, we analyzed citations within experiment-centric sections (e.g., Experiments, Results, Evaluation) as primary anchors. Baseline methods were extracted by analyzing citation contexts and adjacent text, with additional metadata, such as introductory papers and repository links, collected from resources like Papers with Code (PWC) (PWC, [n. d.]) and OpenAlex (Priem et al., 2022). For dataset identification, we used the DataStet (Lopez, 2025) model to automatically extract dataset entities from the experimental sections, validating them through metadata from platforms like Hugging Face (Hug, [n. d.]) and Kaggle (Kag, [n. d.]). This process included entity normalization, alias resolution, and filtering, ensuring high-quality and accurate data for subsequent validation stages.
Figure 5. Illustration of Collective Perception Augmented Retrieval:
Rule-based and LLM-based Filtering.
To further enhance extraction accuracy and effectively mitigate false-positive candidates, we introduced a two-stage filtering strategy. Initially, rule-based filtering mechanisms were formulated to preliminarily evaluate baseline and dataset candidates by leveraging multiple criteria, such as contextual positioning, citation frequency, and naming consistency. Subsequently, entities situated near decision threshold boundaries underwent additional verification through few-shot prompting using large language models (LLMs). Specifically, baseline entities were subjected to LLM-based context-aware evaluation, integrating external metadata, while dataset entities initially underwent metadata-driven normalization and deduplication, followed by high-confidence LLM-based verification combining textual context and metadata. This dual-stage filtering strategy substantially improved overall dataset accuracy and reliability.
4.3. Data Analysis
We conducted an extensive quality assessment and statistical analysis of the finalized scholarly dataset. This comprehensive dataset comprises 108,825 research papers published between 2015 and 2025 across ten leading artificial intelligence conferences, associated with 116,970 distinct baseline methods and 68,316 evaluation datasets. On average, each paper is linked to 5 baselines and 3 datasets, with the completeness ratio of repository links achieving 13%.
A key finding from our analysis is the remarkably high reuse rate of established resources within the community. As illustrated in Figure 4, both baselines and datasets exhibit strong recurring adoption patterns over time. For instance, the baseline reuse rate consistently remains above 70% in recent years, with several periods even exceeding the 90% threshold. This indicates that our dataset successfully captures the core set of canonical methods and benchmarks that form the foundation for experimental comparisons in AI research, rather than just a superficial collection of mentioned names.
These results underscore that our automated extraction and validation pipeline notably surpasses existing methodologies in terms of informational completeness and its ability to map the central, actively-used research components. Consequently, our approach substantially enhances dataset scale, information richness, and breadth of coverage, offering more reliable and comprehensive support for academic recommendation systems and reproducibility initiatives.
5. Method
5.1. Overview
We aim to recommend experimentally suitable baselines and datasets for a target research idea, rather than merely retrieving items with high topical similarity. Most prior approaches (Viswanathan et al., 2023b; Färber and Leisinger, 2021a, b; Beltagy et al., 2019; Qayyum et al., 2025; Goyal et al., 2024; Gu et al., 2022) represent each candidate only from a first–person perspective (its self–description in the introducing paper or repository card). Such views often diverge from how the community actually positions and uses the candidate, omitting crucial third–person signals from citation and usage contexts. In addition, baseline and dataset recommendation are typically treated independently, whereas in real workflows the two are tightly coupled: when selecting datasets we routinely consider those that co–occur with plausible baselines (and vice versa), reflecting a strong baseline–dataset synergy.
Motivated by these observations (Shen and Barabási, 2014), we propose a two–stage recommendation framework. (1) Collective Perception–Augmented Retrieval. We represent each candidate by concatenating its self–description with collective perception, and finetune an embedding model on these representations to improve recall of components. (2) Reasoning–Augmented Reranking. We extract paper-baseline–paper-dataset(paper-dataset–paper-baseline) interaction chains from the literature and train a language model to generate reasoning chains over these interactions to rerank the shortlist.
5.2. Citation Context Augmented Retrieval
This stage builds retrieval representations that jointly encode a candidate’s first–person self–description and third–person collective perception distilled from citation contexts, and then finetunes a dense retriever to maximize recall of baselines and datasets, see fig 5.
Citation Context Extraction
Extracting faithful citation contexts is non-trivial. Mentions of a target $t$ (baseline or dataset) are often ambiguous (acronyms and aliases) and non-indicative of use (generic related-work prose) with multiple targets sometimes packed into one sentence. Naively collecting surrounding sentences therefore injects noise and dilutes the signal about how $t$ is actually positioned and used. To reduce this noise, given a corpus paper $p$, we first locate experiment‐centric sections. For each mention of a target $t$ (baseline or dataset) in $p$, we extract a localized window around the mention consisting of the containing sentence plus its immediate neighbors to capture the stated purpose, setting, and outcome of using $t$. We keep these windows as raw citation contexts without further post-processing. This yields a set $C_{p}(t)=\{c_{1},\dots,c_{m}\}$ for $p$. Because a given target $t$ may be used across many papers, we aggregate all contexts over the corpus as $C(t)=\bigcup_{p}C_{p}(t)$, which provides multiple, potentially diverse, evidence snippets about when and how $t$ is used in practice.
Collective Perception
Our goal is to capture a target’s collective perception. A naive choice is to treat the pooled citation contexts $C(t)$ as this signal. We avoid this because a popular target may accrue thousands of mentions, far beyond the embedding model’s context window. To address these issues, We instantiate a target’s Collective Perception(CP) as a citation context summary. For each target $t$, we use GPT4o (OpenAI, 2024) to synthesize $\mathrm{CP}(t)$ from its pooled citation contexts $C(t)$. The summary answers when and why the baseline or dataset is used, distilling recurring signals about task assumptions, evaluation protocols (metrics, splits, scales), typical configurations, and common co-usage patterns. During synthesis, we de-duplicate paraphrases, discard background or non-experimental mentions, and enforce an evidence-grounded prompt that favors majority-supported observations and avoids speculation.
Target Representation
We adopt early text-level fusion by concatenating the target’s self-description with its collective perception $\mathrm{CP}(t)$:
$x_{t}=\texttt{[Description($t$)]}\ \|\ \texttt{[CP($t$)]}.$
This single-pass encoding lets the encoder’s global self-attention align fine-grained signals across the two segments.
Retriever Finetuning
We finetune a Qwen-based bi-encoder retriever(Qwen3-Embedding-0.6B (Zhang et al., 2025)) with contrastive supervision.
Query and candidate formatting. For a query paper $p$, we form a synopsis $q=\mathrm{Summ}(a(p))$ as in §3. We prepend an instruction prefix to queries only:
$\texttt{Instruct: }\mathcal{T}\ \ \texttt{Query: }q,$
where $\mathcal{T}$ describes the task (e.g., “Given a research idae, retrieve relevant baseline methods or datasets that are most suitable.”). Candidates remain in their native form without an instruction:
$x_{t}=\texttt{[Description($t$)]}\ \|\ \texttt{[CP($t$)]}.$
Let $f_{\theta}(\cdot)$ denote the encoder and $\oplus$ denote string concatenation. We obtain embeddings $h_{q}=f_{\theta}(\texttt{Instruct}\oplus q)$ and $h_{t}=f_{\theta}(x_{t})$, and define the temperature-scaled similarity
$s(q,t)=\tau\,\langle h_{q},h_{t}\rangle,\quad\tau>0.$
Loss. Given a batch $\{(q_{i},t_{i},y_{i})\}_{i=1}^{B}$ with one aligned positive per row (non-diagonal entries serve as in-batch negatives), we compute the similarity matrix $S_{ij}=\tau\,\langle h_{q_{i}},h_{t_{j}}\rangle$ and optimize a diagonal binary cross-entropy with an in-batch contrastive regularizer:
$\mathcal{L}=\underbrace{\frac{1}{B}\sum_{i=1}^{B}\mathrm{BCE}\!\big(S_{ii},\,y_{i}\big)}_{\text{diagonal (main)}}\;+\;\lambda\,\underbrace{\frac{1}{B(B-1)}\sum_{i\neq j}\mathrm{BCE}\!\big(S_{ij},\,0\big)}_{\text{in-batch contrastive}}.$
This objective pulls each query toward its paired positive while pushing it away from other candidates in the batch, without explicit hard-negative mining.
5.3. Reasoning–Augmented Reranking
This stage constructs reasoning chain representations from observed P-B-P-D(P-D-P-B) interactions and finetunes a language model to rerank the shortlist with interpretable, chain-based justifications.
Interaction Chain
[…]
$p\ \xrightarrow{\ \text{dataset}\ }\ d\ \xrightarrow{\ \text{used-by}\ }\ p^{\prime}\ \xrightarrow{\ \text{baseline}\ }\ b,$
which supports baseline recommendation via $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$. Dataset recommendation is symmetric: $p\!\to\!b\!\to\!p^{\prime\prime}\!\to\!d$. For each hop we attach evidential snippets.
Reasoning Chain
To unlock LLM reasoning and enable reasoning reranking based on interaction chains, we use GPT4o (OpenAI, 2024) to generate reasoning Chain trajectories. For each query and each candidate baseline $b$, we enumerate chains of the form $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$ and score each chain by
$m_{D\!\to B}(d,b)\;=\;\bigl|\{\,p^{\prime}\;:\;p^{\prime}\text{ uses }d\text{ and }b\,\}\bigr|$
We retain only the three chains with the largest $m_{D\!\to B}(d,b)$ per candidate baseline and expose these chains and their counts to the LLM as the primary evidence. Symmetrically, for each candidate dataset $d$, we enumerate $p\!\to\!b\!\to\!p^{\prime\prime}\!\to\!d$ and score by
$m_{B\!\to D}(b,d)\;=\;\bigl|\{\,p^{\prime}\;:\;p^{\prime}\text{ uses }b\text{ and }d\,\}\bigr|$
again keeping only the top three chains per candidate dataset. A teacher LLM (GPT-4o) is prompted to reason step-by-step and produce a final ordering, with instructions to rely primarily on the top-three chains and their co-usage metrics, thereby yielding interpretable reranking.
ReRanker Training
To endow the LLM with interaction-chain reasoning, we finetune it as a listwise reranker via supervised finetuning (SFT). Each training instance is a triplet $(Q,R,A)$: $Q$ is the query, $R$ is the chain evidence for each candidate (top-3 interaction chains with co-usage counts $m_{D\!\to B}$ or $m_{B\!\to D}$), and $A$ is the target reasoning chain ending with a normalized decision. The model takes $(Q,R)$ as input and is trained to generate $A$ via token-level negative log-likelihood:
$\theta^{*}\;=\;\arg\min_{\theta}\;\mathbb{E}_{(Q,R,A)\sim\mathcal{D}}\Big[-\sum_{t=1}^{|A|}\log p_{\theta}\!\left(a_{t}\,\middle|\,a_{<t},\,Q,\,R\right)\Big].$
At inference, given $(Q,R)$ the LLM outputs a brief justification and calibrated ranked list, from which we read off the final ranking.
Table 2. Comparison on paper–baseline and paper–dataset recommendation with conservatively estimated Recall@10/20/30 and Hitrate@5/10/15.
Paper–Baseline Paper–Dataset
Method R@10 R@20 R@30 HR@5 HR@10 HR@15 R@10 R@20 R@30 HR@5 HR@10 HR@15
BM25 (Robertson et al., 2009) 0.2080 0.2971 0.3417 0.4239 0.5171 0.5730 0.1334 0.1906 0.2192 0.3143 0.4107 0.4685
Textual-GCL (Qayyum et al., 2025) 0.2516 0.3594 0.4133 0.4457 0.5696 0.6439 0.1168 0.1668 0.1918 0.3273 0.4309 0.4931
SciBERT (Beltagy et al., 2019) 0.2734 0.3905 0.4491 0.5241 0.5965 0.6399 0.1601 0.2287 0.2630 0.3452 0.4674 0.5407
HAtten (Gu et al., 2022) 0.2918 0.4168 0.4793 0.5360 0.6216 0.6730 0.1803 0.2575 0.2961 0.3912 0.4778 0.5298
SymTax (Goyal et al., 2024) 0.2953 0.4218 0.4851 0.5476 0.6453 0.7039 0.2011 0.2873 0.3304 0.4210 0.5125 0.5674
AgentExpt 0.3422 0.4523 0.5036 0.5933 0.6938 0.7387 0.2236 0.3001 0.3533 0.4557 0.5549 0.5968
6. Experiment
6.1. Experimental Setup
Baselines.
We compare our method against five representative families: (i) the lexical matching baseline BM25 (Robertson et al., 2009); (ii) the graph/contrastive approach Textual-GCL (Qayyum et al., 2025); (iii) the scientific-domain encoder SciBERT (Beltagy et al., 2019); (iv) the attention-based model HAtten (Gu et al., 2022); and (v) the structure-aware method SymTax (Goyal et al., 2024). These cover lexical, dense, graph/contrastive, and attention/structure-aware paradigms commonly used for scientific retrieval and recommendation.
Implementation Details.
[…]
6.2. Main Result
On both baseline and dataset recommendation, our method attains the best results in the main comparison. On the baseline side, our retriever improves Recall@20 from 0.4218 (SymTax) to 0.4523 (+0.0305, +7.23%), and the full two-stage model further lifts precision-oriented metrics: HitRate@5 from 0.5476 to 0.5933 (+0.0457, +8.35%) and HitRate@10 from 0.6453 to 0.6938 (+0.0485, +7.52%). Relative to BM25, Recall@20 rises from 0.2971 to 0.4523 (+52.2%). On the dataset side, our retriever increases Recall@20 over SymTax from 0.2873 to 0.3001 (+0.0128, +4.46%), while the full model reaches HitRate@5 0.4557 vs. 0.4210 (+0.0347, +8.24%) and HitRate@10 0.5549 vs. 0.5125 (+0.0424, +8.27%). Compared to BM25, Recall@20 improves from 0.1906 to 0.3001 (+57.5%). As expected, reranking the top-20 does not change Recall@20 but consistently boosts HitRate, especially for datasets (from 0.4250 to 0.4557 at @5, +7.22%; and from 0.5308 to 0.5549 at @10, +4.54%). These gains align with our design: collective perception–enhanced retrieval raises recall, and interaction chain–guided reasoning improves precision by selecting components that are experimentally suitable rather than merely topically similar.
6.3. Ablation Study
Ablation Study of Retriever.
Removing collective perception (CP) causes the largest drop. On the baseline side, adding CP over “w/o CP” yields +0.0807 Recall@20 (+21.72%), +0.0998 HitRate@5 (+21.15%), and +0.1066 HitRate@10 (+18.73%); adding self-description on top of CP still contributes (+0.0326 Recall@20, +7.77%; +0.0368 @5, +6.88%; +0.0546 @10, +8.79%). The dataset side shows the same pattern: CP accounts for the majority of the gains (+0.0508 Recall@20, +20.38%; +0.0958 @5, +29.10%; +0.1146 @10, +27.53%), while self-description brings steady, smaller improvements (+0.0292 Recall@20, +10.78%; +0.0234 @5, +5.83%; +0.0330 @10, +6.63%). This confirms that community citation contexts supply decisive third-person signals about how components are used, beyond first-person descriptions.
Ablation Study of Reranker.
Both interaction chains and supervised fine-tuning (SFT) are crucial. On the baseline side, removing interaction chains reduces HitRate@5 by 0.0688 (-13.12% relative) and HitRate@10 by 0.0583 (-9.17% relative); removing SFT yields an even larger drop (-0.1649 @5, -38.49% relative; -0.1126 @10, 19.37% relative). The dataset side mirrors this: -0.0431 @5 (-10.45% relative) and -0.0502 @10 (-9.95% relative) without chains; and -0.0916 @5 (-25.16% relative) and -0.1301 @10 (-30.63% relative) without SFT. Overall, interaction chains provide high-precision evidential priors about co-usage, and SFT teaches the model to exploit these priors to make calibrated, interpretable decisions—together delivering the observed precision gains.
Ablation Study of Recommendation Pipeline.
To assess the impact of the reranking module, we performed an ablation by removing the reranker (denoted w/o Reranker) and comparing its performance to the full model. As shown in Table 5, removing the reranker caused a notable drop in HitRate@5 and HitRate@10 for both the baseline and dataset recommendation tasks. This performance loss shows that while the retriever identifies candidate items, it cannot effectively prioritize the most relevant ones. In contrast, the reranking component, utilizing interaction chains and contextual reasoning, refines rankings and improves both ranking quality and recommendation precision. These results highlight the importance of integrating retrieval and reranking, demonstrating that the reranker significantly enhances the recommendation pipeline’s effectiveness.
Method R@20 HR@5 HR@10
Baseline w/o Collective Perception 0.3716 0.4719 0.5691
w/o Self-Description 0.4197 0.5349 0.6211
AgentExpt(full) 0.4523 0.5717 0.6757
Dataset w/o Collective Perception 0.2493 0.3292 0.4162
w/o Self-Description 0.2709 0.4016 0.4978
AgentExpt(full) 0.3001 0.4250 0.5308
Table 3. Ablation study of retriever
Method Hitrate@5 Hitrate@10
Baseline w/o Reasoning Argumented 0.4284 0.5812
w/o Interaction Chain 0.5245 0.6355
AgentExpt(full) 0.5933 0.6938
Dataset w/o Reasoning Argumented 0.3641 0.4248
w/o Interaction Chain 0.4126 0.5047
AgentExpt(full) 0.4557 0.5549
Table 4. Ablation study of reranker
Method Hitrate@5 Hitrate@10
Baseline w/o Reranker 0.5717 0.6757
AgentExpt(full) 0.5933 0.6938
Dataset w/o Reranker 0.4250 0.5308
AgentExpt(full) 0.4557 0.5549
Table 5. Ablation study of recommendation pipeline
7. Conclusion
We proposed a unified framework for recommending baselines and datasets. On the data side, we built a large-scale resource linking papers to components across ten AI venues. On the method side, we introduced collective perception and fine-tuned a dense retriever to improve recall. We also developed an interaction-chain-guided reranker to enhance reasoning and interpretability. Experiments and case studies show consistent improvements over baselines, highlighting the role of community signals and chain reasoning.